300g of water at 25C is added to 100g 0f ice at 0C. The final temperature of mixture is 
  1. -5/2C
  2. -5/3C
  3. -5C
  4. 0C
The answer is 0○ but I got -5/4 using Q1 + Q2=Q3 formula PLEASE HELP!!!!

Grade:12th pass

3 Answers

Rinki jain
6 years ago
1 st find heat content of water by Q=mst And latent heat req. To melt ice by Q=ml ... U will see that water is not able to melt the entire ice.... Therefore some ice will remain and temp. Will be 0 Degree Celsius
Yash Chourasiya
3 years ago
Dear Student

The heat required for 100g of ice at 0C to change into
Water at0C = mL = 100×80×4.2 = 33,600J…(i)
The heat released by 300g of water at 25C to change its
temperature to 0C = mcΔT = 300×4.2×25 = 31,500J….(ii)
Since the energy in eq.(ii) is less than of eq(i)
therefore the final temperature will be 0C

I hope this answer will help you.
Thanks & Regards
Yash Chourasiya
ssssssssss
2 years ago
dhuodshaoshdoaidwapoisjpsdwdoahdsjdsjksdhpaikdpoadWDWDWDdadsadawfqdsafffwwfsdadsdQAdqADdqsqASqawdaafwaf